# Bayesian Quantile Regression with Subset Selection: A Posterior Summarization Perspective

### v3

@@ 2.1 Point Estimation for Linear Quantiles
with zeros for indices $j\notin S$, where $\hat{Q}_{\tau}(\boldsymbol{x}_{i})=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\{Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})\}$, $\boldsymbol{\hat{Q}}_{\tau}(\boldsymbol{X})=\{\hat{Q}_{\tau}(\boldsymbol{x}_{1}),\dots,\hat{Q}_{\tau}(\boldsymbol{x}_{n})\}^{\intercal}$, and ${\boldsymbol{X}}_{S}$ the $n\times\lvert S\rvert$ matrix of active covariates for subset $S$.
Thus, the optimal linear action under the quantile-focused squared error loss is the least squares solution with response vector $\boldsymbol{\hat{Q}}_{\tau}({\boldsymbol{X}})$ and covariate submatrix ${\boldsymbol{X}}_{S}$. The result is a linear point estimate of the $\tau$th conditional quantile of $Y\mid\boldsymbol{x}$, which provides the magnitude and direction of the relationship between each covariate and a specific percentile of the response. These coefficients may also be used for quantile-specific prediction.
In general, the posterior expected conditional quantiles $\hat{Q}_{\tau}(\boldsymbol{x}_{i})=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\{Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})\}$ under $\mathcal{M}$ are not available in closed-form. However, Monte Carlo approximations are easily obtained using posterior samples $\{\boldsymbol{\theta}^{m}\}_{m=1}^{M}\sim p(\boldsymbol{\theta}\mid\boldsymbol{y})$ via $E_{\boldsymbol{\theta}\mid\boldsymbol{y}}[Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})]\approx M^{-1}\sum_{m=1}^{M}Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta}^{m})$.
@@
While Corollaries 1 and 2 confirm reasonable behavior for the optimal action (6) under linear regression models, we emphasize the utility of Lemma 2.1 for quantile regression estimation under any Bayesian regression model. In particular, the analyst can prioritize curation of $\mathcal{M}$ to capture complex features of the entire conditional distribution $Y\mid\boldsymbol{x}$, including nonlinearity, skewness, and heteroscedasticity, while (8) provides the optimal linear approximation of the model-based quantiles under the loss function (7).
@@ 3.1 Subset Search with Decision Analysis
which resembles (14) but instead inherits a posterior distribution via $p(\boldsymbol{\theta}\mid\boldsymbol{y})$ under model $\mathcal{M}$. Thus, (15) enables uncertainty quantification for the performance of each subset. The evaluation criteria in (14) and (15) are used in conjunction to search and subsequently filter to the “near-optimal” or acceptable family of subsets (Kowal, 2021), specific to each quantile $\tau$.
@@ 3.3 Quantifying Variable Importance
A principal advantage of curating the quantile-specific acceptable family (17) is that it contains more information than any single (“best”) subset. In the common applied setting with moderate $p$, correlated covariates, or weak signals, there are typically many subsets that offer similar accuracy. Thus, we de-emphasize the selection of a single ”best” subset and instead provide more comprehensive summaries of the quantile-specific acceptable family.
@@ 4.4 Selection
5 Quantile Estimation, Inference, and Selection for Educational Outcomes
We now proceed to the quantile regression analysis of the North Carolina data, aiming to uncover the social stressors, environmental exposures, and other factors that differentially impact low-, medium-, and high-achieving students. We augment the covariates in Table 1 with interactions between mother’s race (mRace) and each of blood lead level (blood_lead), neighborhood deprivation (NDI), and racial residential isolation (RI). Crucially, these interactions allow us to assess whether the impacts of these social and environmental exposures differ among quantiles and between race groups.

### v4

@@ 2.1 Point Estimation for Linear Quantiles
with zeros for indices $j\notin S$, where ${\boldsymbol{X}}_{S}$ is the $n\times\lvert S\rvert$ matrix of active covariates for subset $S$, $\boldsymbol{\hat{Q}}_{\tau}(\boldsymbol{X})=\{\hat{Q}_{\tau}(\boldsymbol{x}_{1}),\dots,\hat{Q}_{\tau}(\boldsymbol{x}_{n})\}^{\intercal}$, and $\hat{Q}_{\tau}(\boldsymbol{x}_{i})=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\{Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})\}$.
The optimal linear action under the quantile-focused squared error loss is the least squares solution with response vector $\boldsymbol{\hat{Q}}_{\tau}({\boldsymbol{X}})$ and covariate submatrix ${\boldsymbol{X}}_{S}$. The result is a linear point estimate of the $\tau$th conditional quantile of $Y\mid\boldsymbol{x}$, which provides the magnitude and direction of the relationship between each covariate and a specific percentile of the response. These coefficients may also be used for quantile-specific prediction.
In general, the posterior expected conditional quantiles $\hat{Q}_{\tau}(\boldsymbol{x}_{i})=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\{Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})\}$ under $\mathcal{M}$ are not available in closed-form. Monte Carlo approximations are easily obtained via $E_{\boldsymbol{\theta}\mid\boldsymbol{y}}[Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})]\approx M^{-1}\sum_{m=1}^{M}Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta}^{m})$ where $\{\boldsymbol{\theta}^{m}\}_{m=1}^{M}\sim p(\boldsymbol{\theta}\mid\boldsymbol{y})$.
@@
While Corollaries 1 and 2 confirm reasonable behavior for the optimal action (6) under linear regression models, we emphasize the utility of Lemma 2.1 for quantile regression estimation under any Bayesian regression model. In particular, the analyst can prioritize curation of $\mathcal{M}$ to capture complex features of the entire conditional distribution $Y\mid\boldsymbol{x}$, including nonlinearity, skewness, and heteroscedasticity, while (8) provides the optimal linear approximation of the model-based quantiles under the quantile-focused loss (7).
@@ 3.1 Subset Search with Decision Analysis
which resembles (14) but instead inherits a posterior distribution via $p(\boldsymbol{\theta}\mid\boldsymbol{y})$ under model $\mathcal{M}$. Thus, (15) enables uncertainty quantification for the performance of each subset. The evaluation criteria in (14) and (15) are used in conjunction to search and subsequently filter to the ”near-optimal” or acceptable family of subsets (Kowal, 2021), specific to each quantile $\tau$.
@@ 3.3 Quantifying Variable Importance
A principal advantage of curating the quantile-specific acceptable family (17) is that it contains more information than any single (”best”) subset. In the common applied setting with moderate $p$, correlated covariates, or weak signals, there are typically many subsets that offer similar accuracy. Thus, we de-emphasize the selection of a single ”best” subset and instead provide more comprehensive summaries of the quantile-specific acceptable family.
@@ 4.4 Selection
5 Social Stressors, Environmental Exposures, and Childhood Educational Outcomes
Childhood educational outcomes are affected by social stressors (e.g., poverty, structural racism, high unemployment) and environmental exposures (e.g., poor air quality, lead exposure), which often cumulate in the same communities (Miranda et al., 2007; Bravo et al., 2024). Thus, there is urgency to quantify and characterize these effects and initiate well-informed policy interventions, especially given the strong links between educational attainment and adult health. However, far less is known about the differential impacts of social stressors and environmental exposures on low-, medium-, and high-achieving students. Such information is critical for understanding the factors that shape educational outcomes and for recommending policy interventions, especially to support at-risk students. Thus, quantile regression is an informative tool—especially with the ability to quantify effect directions and magnitudes, measure and report uncertainty, and provide quantile-specific subset selection.
[…]
Birth information
mEdu
Mother’s education group at the time of birth
(No high school diploma, High school diploma,
College diploma)
mRace
Mother’s race/ethnicity group (Non-Hispanic (NH) White,
NH Black)
BWTpct Birthweight percentile
mAge Mother’s age at the time of birth
Male Male infant? (1 = Yes)
Smoker Mother smoked? (1 = Yes)
NotMarried Not married at time of birth (1 = Yes)
NOPNC
Mother received pre-natal care before birth?
(1 = No prenatal care)
Weeks_Gestation Gestational period (in weeks)
Education/End-of-grade (EoG) test information
Reading_Score
Standardized score for the (chronologically first) 4th EoG
reading test
Blood lead surveillance
Blood_lead Blood lead level (micrograms per deciliter)
Social/Economic status
EconDisadvantage
Participation in Child Nutrition Lunch Program? (1 = Yes)
NDI Neighborhood Deprivation Index, at time of EoG test
RI Racial residential isolation, at time of EoG test
We augment the covariates in Table 3 with interactions between mother’s race (mRace) and each of blood lead level (blood_lead), neighborhood deprivation (NDI), and racial residential isolation (RI). Crucially, these interactions allow us to assess whether the (possibly heterogeneous) effects of environmental exposures and social stressors on educational outcomes also differ between race groups.